H-80.2 B)69.3 < μ < 79.7 C) 70.5 < μ < 79.5 12. In a time-use study, 20 randomly selected managers were found to spend a mean of 2.4 hours each day on paperwork. The standard deviation of the 20 scores is s 1.3 D) 67.7 < μ < 84.7 ours. Construct the 95% confidence interval for the population standard deviation σ (Use y -distribution for analysis). Provide estimates and circle one of the multiple choice answers. (6 points) A) 1.05 < σ < 1.81 B) 0.99 < σ <1.90 C)0.93 < σ <1.74 D) 1.17 < σ < 2.03 2·The nighttime cold medicine Dozenol bears a label aminophen in each fluid ounce of the drug d 36 1-ounce samples and found andard deviation is 21 2.hh36d denlain
Answer #1

solution:- option B) 0.99 < σ < 1.90

given that n = 20 ,

mean = 2.4 , standard deviation s = 1.3

95% confidence and with df = 20 - 1 = 19

=>X^2 distribution

=> X^2 R = 32.85

=> X^2 L = 8.907

formula

=> sqrt((n-1)* s^2 /X^2 R) < σ < sqrt((n-1)* s^2 /X^2 L)

=> sqrt(20-1)* 1.3^2 /32.85) < σ < sqrt(20-1)* 1.3^2 /8.907)

=> sqrt(19*1.69/32.85) < σ < sqrt(19* 1.69 /8.907)

=> 0.99 < σ < 1.90